# Generalized local jocabians and commutative group stacks

### v2

@@ 4.1. Affine homology
$\mathbf{Map}_{\mathbf{AbChAff}_{k}}(Gr_{d}\mathbb{H}^{u}(X),E))\simeq\oplus_{x\in X^{(d)}}\mathbb{H}^{*}_{x}(X,E).$
Because affine homology type of finite presentation generate the whole $\infty$-category $\widehat{\mathbf{AbChAff}_{k}}$ by limits, we deduce that we have we have a general formula in terms of local affine homology
$Gr_{d}\mathbb{H}^{u}(X)\simeq\prod_{x\in X^{(d)}}\mathbb{H}^{u}_{x}(X).$
@@ 4.2. Affine homology as a dg-module
The above corollary implies that for a small stack $X$ its affine homology $\mathbb{H}^{u}(X/k)$ is characterized by $\mathbb{R}\Gamma(X,\mathbb{G}_{a})$, as a dg-module over $C_{\mathbb{G}_{a}}$. The dg-algebra $C_{\mathbb{G}_{a}}$ is however quite complicated in general. It simplifies when $\mathbb{G}_{a}$ is restricted to sites of perfect schemes (see [Bre81]). In particular, things tend to simplify when the base ring $k$ is perfect (or at least perfectoid). Another simplification consists of using the group of big Witt vectors $\mathbb{W}$ instead of the additive group $\mathbb{G}_{a}$, as for instance $Ext^{1}(\mathbb{W},\mathbb{W})=0$, and again higher ext vanish when computed over the site of perfect schemes. Moreover, $End(\mathbb{W})$ is the usual Dieudonné ring so dg-modules over $\mathbb{R}End(\mathbb{W})$ are closed related to usual Dieudonné modules. However, $\mathbb{W}$ is not a cocompact object in $\widehat{\mathbf{AbChAff}}_{k}$, and therefore the corresponding adjunction with dg-modules is not as well behaved and relates dg-modules over $\mathbb{R}End(\mathbb{W})$ and certain pro-objects inside $\mathbf{AbChAff}_{k}$.

### v3

@@ 4.1. Affine homology
$\mathbf{Map}_{\mathbf{AbChAff}_{k}}(Gr_{d}\mathbb{H}^{u}(X),E))\simeq\oplus_{x\in X^{(d)}}\mathbb{H}_{x}(X,E).$
Because affine homology type of finite presentation generate the whole $\infty$-category $\widehat{\mathbf{AbChAff}_{k}}$ by limits, we deduce that we have a general formula in terms of local affine homology
$Gr_{d}\mathbb{H}^{u}(X)\simeq\prod_{x\in X^{(d)}}\mathbb{H}^{u}_{x}(X),$
where each factor $\mathbb{H}^{u}_{x}(X)$ is defined as the cofiber, taken in $\widehat{\mathbf{AbChAff}_{k}}$ of the natural morphism $\mathbb{H}^{u}(X_{x}-\{x\})\to\mathbb{H}^{u}(X_{x})$.
@@ 4.2. Affine homology as a dg-module
The above corollary implies that for a small stack $X$ its affine homology $\mathbb{H}^{u}(X/k)$ is characterized by $\mathbb{R}\Gamma(X,\mathbb{G}_{a})$, as a dg-module over $C_{\mathbb{G}_{a}}$. The dg-algebra $C_{\mathbb{G}_{a}}$ is however quite complicated in general. It simplifies when $\mathbb{G}_{a}$ is restricted to sites of perfect schemes (see [6]). In particular, things tend to simplify when the base ring $k$ is perfect (or at least perfectoid). Another simplification consists of using the group of big Witt vectors $\mathbb{W}$ instead of the additive group $\mathbb{G}_{a}$, as for instance $Ext^{1}(\mathbb{W},\mathbb{W})=0$, and again higher ext vanish when computed over the site of perfect schemes. Moreover, $End(\mathbb{W})$ is the usual Dieudonné ring so dg-modules over $\mathbb{R}End(\mathbb{W})$ are closed related to usual Dieudonné modules. However, $\mathbb{W}$ is not a cocompact object in $\widehat{\mathbf{AbChAff}}_{k}$, and therefore the corresponding adjunction with dg-modules is not as well behaved and relates dg-modules over $\mathbb{R}End(\mathbb{W})$ and certain pro-objects inside $\mathbf{AbChAff}_{k}$.